# Reverse Diffusion Sequential Monte Carlo Samplers

### v1

@@ 4 Related works
Training-free diffusion-based samplers.
Huang et al. [8] develops a sampling algorithm based on the reverse SDE of an Ornstein-Uhlenbeck process, using MCMC to estimate the score function. This method, in contrast to RDSMC, initializes the reverse process at an earlier time $\tau_{T}<1$, where initial samples are obtained via a nested Langevin procedure. Grenioux et al. [9] considers a similar stochastic localization scheme. They further adapt the time discretization of the SDE to its signal-to-noise ratio (SNR) profile, and provide guidance on choosing the initial sampling time $\tau_{T}$ given certain target conditions.
While Huang et al. [8] provides an upper bound on the sample complexity, their practical performance can be hindered by finite time discretization and score approximation biases. RDSMC wraps similar diffusion-based samplers as a proposal within a principled SMC framework. By introducing novel intermediate targets, we systematically correct for discretization and score estimation errors and provide an unbiased estimate of the normalization constant.
Diffusion-based neural samplers.
A separate line of work on neural samplers trains diffusion model for sampling. Variational approaches achieve this goal by minimizing the divergence between forward and reverse processes [13, 14, 16, 17, 15], while Akhound-Sadegh et al. [11], OuYang et al. [12] exploit analytic expressions of the score function or a related energy function. Similar to their training-free counterparts, these methods are prone to discretization and approximation errors.
Closest to our work are Phillips et al. [10], Chen et al. [18] that also use SMC to reduce errors in neural samplers. Unlike our method, their proposals require training. Phillips et al. [10] construct intermediate targets using a learned neural network approximation, while ours are training-free and reflect the true target marginals at each SMC iteration. Chen et al. [18] define the extended target along a geometric annealing path, while we derive ours from the forward process of diffusion models.
Recently, He et al. [20] show that neural samplers typically incur heavy computational overhead and rely on specific network conditioning, compared to classical sampling methods. For this reason, we limit our empirical comparison to training-free samplers. Nonetheless, an interesting direction for future work is to combine our method with training-based approaches to further enhance performance.
@@
Beyond classical sampling tasks, SMC has been applied to conditional generation for pre-trained diffusion models [29, 30, 31, 32]. While these works also combine SMC and diffusion dynamics, a key distinction is the target distribution: they sample from an existing diffusion model prior tilted with a reward function, whereas our setting involves sampling from any (unnormalized) target distribution. As a result, the way diffusion models are used is different, and we require distinct designs for proposals and intermediate targets in SMC.
@@ 5 Experiments
We evaluate RDSMC on a range of synthetic and real-world target distributions, comparing it to SMC [25], AIS [24], SMS [33], RDMC [8], and SLIPS [9]. AIS and SMC operate over a series of geometric interpolations between the target and a Gaussian proposal with MCMC transitions. SMS samples from the target by sequentially denoising equally noised measurements using a Langevin procedure combined with estimated scores. RDMC and SLIPS are described in §˜4.
RDSMC uses a variance-preserving diffusion schedule with annealed importance sampling for score estimation. To evaluate the effectiveness of intermediate target distributions and importance of the resampling mechanism, we include two ablations: (i) RDSMC (Proposal), which samples directly from the reverse diffusion proposal in Eq.˜8 without any weighting or resampling, and (ii) RDSMC (IS), which applies a final IS correction to samples from RDSMC (Proposal).
Baseline methods follow the implementation of Grenioux et al. [9]. Notably, the information about the target variance (or an estimation) is provided to guide the initialization of the baselines SMC, AIS, SMS and SLIPS. In contrast, our method does not make use of this information.
Unless otherwise specified, we use $T=100$ discretization steps for RDSMC and its variants, and $T=1,024$ steps for other methods. We generate $N=4,096$ final samples for all methods, and tune their hyperparameters assuming access to either a validation dataset or an oracle metric. All experiment details are included in Appendix˜D.
@@ 5.1 Bi-modal gaussian mixtures
(b) Estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$ and log-normalization constant $\log Z$ versus dimension $d$. Results are averaged over 5 seeds with error bars showing one standard error. Bias increases with $d$ for all methods, while RDSMC consistently outperforms the baselines. SMS, RDMC, and SLIPS results are shown in §˜D.2 and exhibit consistently higher bias.
@@
We first study bi-modal Gaussian mixtures with an imbalanced weight ratio of $w_{1}/(w_{1}+w_{2})=0.1$ for varying dimensions $d$. The estimated ratio is obtained by assigning samples to the mode with the highest posterior probability. For each method we report the result with the lowest estimation bias.
@@
Figure˜1(b) shows the estimation bias of the weight ratio (left), and that of the log-normalization constant $\log Z$ (right). RDSMC consistently outperforms AIS and SMC in both metrics across dimensions. RDSMC (Proposal) exhibits consistently high weight ratio estimation bias. While RDSMC (IS) reduces this bias, it still underperforms the full RDSMC procedure, highlighting the effectiveness of intermediate resampling guided by our intermediate target distributions.
@@
Results for SMS, RDMC, and SLIPS are included in §˜D.2, which show higher weight ratio estimation bias compared to RDSMC. Additionally, they do not provide estimates for $\log Z$.
@@ 5.2 Rings and Funnel distributions
Table 1: Synthetic target results (mean $\pm$ standard error over 5 seeds). Bold indicates 95% confidence interval overlap with that of the best average result. RDSMC consistently yields the lowest $\log Z$ estimation bias. On Rings, RDSMC(Proposal) has the lowest radius TVD, followed by AIS, SMC, and RDSMC. On Funnel, SLIPS has the lowest sliced KSD, followed by AIS, SMC, and RDSMC.
We present results on two additional synthetic targets. Rings, introduced by Grenioux et al. [9], is a 2-dimensional distribution constructed via an inverse polar parameterization: the angular component is uniformly distributed, while the radius component follows a 4-mode Gaussian mixture. We assess sample quality using total variational distance in the radius component (radius TVD). Funnel, proposed by Neal [34], is a 10-dimensional “funnel"-shaped distribution. We use sliced Kolmogorov-Smirnov distance (sliced KSD) to evaluate sample quality following Grenioux et al. [9]. Both targets admit tractable normalization constants, so we also report the estimation bias for methods that compute it. For each method, we show results with the highest sample quality on a validation dataset.
As shown in Table˜1, RDSMC achieves lower or comparable estimation bias of the log-normalization constant relative to AIS and SMC on both targets. On the simpler Rings target, RDSMC (Proposal) has the lowest radius TVD, where AIS, SMC and RDSMC closely match. On the more challenging Funnel target, SLIPS achieves the lowest sliced KSD; while RDSMC performs slightly worse, it greatly outperforms its IS and Proposal variants. Additional results are provided in §˜D.3.
@@ 5.3 Bayesian logistic regression
Finally, we evaluate inference performance on Bayesian logistic regression models using four datasets. Credit and Cancer [35] involve predicting credit risk and breast cancer recurrence, while Ionosphere [36] and Sonar [37] focus on classifying radar and sonar signals, respectively. The inference is performed on $60\%$ of each dataset, leaving 20% for validation and 20% for testing. Each method is tuned using the validation log-likelihood estimate. We report the test log-likelihood in Table˜2.
@@ 6 Discussion
As a nested SMC procedure, our proposed method inherently introduces auxiliary variance. Such variance may be mitigated by enhancing autocorrelation within the Monte Carlo estimation step, in the spirit of [38]. In high-dimensional Gaussian mixture experiments, we observe a degree of mode collapse, an issue also seen in other samplers. Potential remedies include adopting partial resampling strategies [39], or incorporating target information into our procedure. Our method also relies on hyperparameter tuning guided by oracle metrics, which aligns with common practice in the related literature. While heuristics aid in this process, developing a more automated strategy remains an important future direction. Finally, to further improve performance, we can incorporate SNR-adapted discretization schemes and informative initializations, as explored by Grenioux et al. [9], as well as combine our approach with diffusion-based neural samplers.
@@ A.1 Diffusion models
In Variance-Preserving (VP) schedule [6], the coefficients of the forward SDE in Eq.˜1 are given by
$f(\tau)=-\frac{1}{2}b(\tau),\quad g(\tau)=\sqrt{b(\tau)},$ (16)
where $b(\tau)$ is a time-dependent function
$b(\tau)=b_{\min}+\tau(b_{\max}-b_{\min}),$
@@
By Ito’s calculus (see e.g. Särkkä and Solin [40]), one can show that the transition kernel $p_{\tau}(x(\tau)\mid x_{0})=\mathcal{N}(x(\tau)\mid\alpha(\tau)x_{0},\sigma(\tau)^{2})$ with
$\alpha(\tau)=\exp\left(-\frac{1}{2}\int_{0}^{\tau}b(s)ds\right),\quad\sigma(\tau)^{2}=1-\exp\left(-\int_{0}^{\tau}b(s)ds\right).$
@@
In Variance-Exploding (VE) schedule [7], the forward SDE coefficients in Eq.˜1 are given by
$f(\tau)=0,\qquad g(\tau)=\sqrt{\frac{\mathrm{d}\sigma^{2}(\tau)}{\mathrm{d}\tau}},$
where $\sigma(\tau)$ is a non-negative and monotone increasing function with $\sigma(0)=0$. Commonly used forms of $\sigma(\tau)$ include polynomial functions, such as linear, quadratic or exponential schedules. When $\sigma(1)$ is large, the reference distribution can be effectively approximated as $p_{1}(x)\approx\mathcal{N}(x\mid 0,\sigma(1)^{2})$.
Similarly to the VP case, one can derive the following parameters of the Gaussian transition kernel $p_{\tau}(x(\tau)\mid x_{0})$ with $\alpha(\tau)=1$ and $\sigma(\tau)$ some non-negative and monotone function.
A.2 Diffusion-based samplers and score identities
@@
TSI utilizes known structural properties of the target density, such as estimated gradients, facilitating potentially more accurate and lower-variance score estimations at low-noise settings. Formally, the score function of the noisy target distribution $p_{\tau}(x)$ under the TSI framework satisfies
$\nabla_{x}\log p_{\tau}(x)$ $=\frac{1}{\alpha(\tau)}\int\nabla_{x}\log p_{\tau}(x_{0})p_{\tau}(x_{0}\mid x)dx_{0}.$ (17)
TSI and its related alternative have been used in several recent studies to improve score estimation accuracy [41, 10, 11, 42].
@@
To balance the advantages of DSI and TSI across different noise regimes, MSI combines both identities through a convex interpolation. Specifically, MSI favors TSI in the low-noise regime and DSI in the high-noise regime, leveraging both strengths. Formally, MSI is a convex combination of TSI and DSI with coefficients $\frac{\alpha(\tau)^{2}}{\alpha(\tau)^{2}+\sigma(\tau)^{2}}$ and $\frac{\sigma(\tau)^{2}}{\alpha(\tau)^{2}+\sigma(\tau)^{2}}$:
$\nabla_{x}\log p_{\tau}(x)$ $=\frac{1}{\alpha(\tau)^{2}+\sigma(\tau)^{2}}\int\Big{(}\alpha(\tau)\big{(}x_{0}+\nabla_{x}\log p_{\tau}(x_{0})\big{)}-x(\tau)\Big{)}p_{\tau}(x_{0}\mid x)dx_{0}.$ (18)
He et al. [43] leverage MSI to train an implicit generative model that produces samples in a single step, demonstrating improved sample quality and efficiency.
@@ Appendix B Method details
In this section, we present detailed descriptions of the score estimation strategies, resampling schemes, and numerical techniques used to improve our method. We also discuss the associated computational complexity.
B.1 Score estimation
In this subsection, we present several score estimators, including IS, AIS and SMC. AIS and SMC can be used to improve score estimates for complex targets.
B.1.1 IS-based score estimator
@@
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
1 Sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}:={\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(0:M)}}$ i.i.d. from some proposal $q$;
2 Compute importance weights: $w^{(m)}\leftarrow\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}},\sigma_{t}^{2}I)}{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}$ for $m=1,\cdots,M$;
3 Compute score estimate: $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}}$ ;
Compute marginal estimate: $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}w^{(m)}$.
Algorithm 2 Importance sampler of $p(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$
The IS-based estimator is summarized in Algorithm˜2. Specifically, we relate the score function to the posterior mean via Tweedie’s formula [44]:
$\nabla_{x_{t}}\log p(x_{t})=\frac{\alpha_{t}\mathbb{E}[{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}]-x_{t}}{\sigma_{t}^{2}}.$
This connection reduces the problem of estimating the score function to estimating the posterior mean $\mathbb{E}[{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}]$. Importance sampling is used to approximate the posterior distribution. We sample from a proposal distribution $q$ and reweight the samples $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} according to their likelihood under the intermediate target distributions. We also derive the marginal estimate as a byproduct.
B.1.2 AIS and SMC-based score estimator

### v2

@@ 4 Related works
Diffusion MC samplers.
huang2023reverse develop the RDMC sampler based on the reverse diffusion process using Langevin-style MCMC to estimate the score function. They initialize the sampler at some $\tau_{T}<1$ via a nested Langevin procedure. grenioux2024stochastic propose a similar SLIPS sampler with a signal-to-noise-ratio (SNR)-adjusted discretization scheme, and provide further guidance on choosing the initial sampling time $\tau_{T}$ under suitable conditions. In contrast, we use AIS-based score estimators to enable marginal estimation and start sampling from the base distribution at $\tau_{T}=1$.
Alternatively, he2024zeroth explore rejection sampling for score estimation, improving the sampling efficiency in low-dimensional settings, which may be incorporated into our framework as well.
While these methods provide certain theoretical guarantees, we offer an orthogonal means to improve their accuracy by increasing the number of particles. RDSMC can be viewed as a generic SMC wrapper around these diffusion-MC samplers (given suitable score estimators), enabling consistent sampling and unbiased estimation of the normalization constant.
Training-based diffusion samplers
In addition to developing MC estimates of the score function, another line of work trains a neural network directly using unnormalized target information. Variational approaches achieve this goal by minimizing the divergence between forward and reverse processes [zhang2021path, berner2022optimal, vargas2023denoising, vargas2023transport, richter2023improved], while akhound2024iterated, ouyang2024bnem exploit various identities of the score function or a related energy function. Similar to their training-free MC-based counterparts, these methods are prone to discretization and score approximation errors.
Closest to our work is phillips2024particle that also use SMC to correct proposal bias. Their method requires training neural networks, while ours is training-free. Moreover, their neural network-based intermediate targets incur approximation errors except at the last step, while ours reflect the true target marginals at each SMC iteration. See an empirical comparison in §˜E.5.3.
he2025no show that many training-based methods incur heavy computational overhead compared to classical sampling methods and rely on special network preconditioning. For this reason, we limit our empirical comparison to training-free samplers. Nonetheless, an interesting direction for future work is to combine our method with training-based approaches to further enhance performance.
@@
Beyond classical sampling tasks, SMC has been applied to conditional generation for pre-trained diffusion models [trippe2022diffusion, wu2023practical, cardoso2023monte, singhal2025general]. While these works also combine SMC and diffusion dynamics, a key distinction is the target distribution: they sample from an existing diffusion model prior tilted with a reward function, whereas our setting involves sampling from any (unnormalized) target distribution. As a result, the way diffusion models are used is different, and we require distinct designs for proposals and intermediate targets in SMC.
Nevertheless, RDSMC uses the estimated marginal densities in Eq.˜11 as twisting functions to improve SMC sampling efficiency. This strategy is conceptually is similar to that of wu2023practical, though in a different setting; see Appendix˜D for further discussion.
@@ 5 Experiments
We evaluate RDSMC on a range of synthetic and real-world target distributions, comparing it to SMC [del2006sequential], AIS [neal2001annealed], SMS [saremi2023chain], RDMC [huang2023reverse], and SLIPS [grenioux2024stochastic]. AIS and SMC operate over a series of geometric interpolations between the target and a Gaussian proposal with MCMC transitions. SMS samples from the target by sequentially denoising equally noised measurements using a Langevin procedure combined with estimated scores. RDMC and SLIPS are described in §˜4.
RDSMC uses a variance-preserving diffusion schedule. To evaluate the effectiveness of our intermediate targets and the importance of resampling, we include two ablations: (i) RDSMC (Proposal), which samples directly from the reverse diffusion proposal in Eq.˜8 without any weighting or resampling, and (ii) RDSMC (IS), which applies a final IS correction to samples from RDSMC (Proposal).
Baseline methods follow the implementation of grenioux2024stochastic. Notably, the information about the target variance (or an estimation) is provided to guide the initialization of the baselines SMC, AIS, SMS and SLIPS. In contrast, our method does not make use of this extra information.
Unless otherwise specified, we use $T=100$ discretization steps for RDSMC and its variants, and $T=1,024$ steps for other methods. We generate $N=4,096$ final samples for all methods, and tune their hyperparameters assuming access to either a validation dataset or an oracle metric.
We provide ablation studies controlling for the discretization steps, the total running time and comparable hyperparameter settings in §˜E.5, where we observe RDSMC still remains competitive and, in some cases, superior. All experiment details are included in Appendix˜E.
@@ 5.1 Bi-modal gaussian mixtures
(b) Estimation bias of weight ratio $w_{1}/(w_{1}+w_{2})$ and log-normalization constant $\log Z$ versus dimension $d$. Results are averaged over 5 seeds with error bars showing one standard error. Note that SMS, RDMC, and SLIPS do not provide estimate of $\log Z$. For both metrics, the bias increases with $d$ for all methods, while RDSMC outperforms the baselines for most cases.
@@
We first study bi-modal Gaussian mixtures with an imbalanced weight ratio of $w_{1}/(w_{1}+w_{2})=0.1$ for varying dimensions $d$. The estimated ratio is obtained by assigning samples to the mode with the highest posterior probability. For each method we select the hyperparameters based on the lowest estimation bias of the weight ratio.
@@
Figure˜1(b) shows the estimation bias of the weight ratio (left), and that of the log-normalization constant $\log Z$ (right). Note that only RDSMC, RDSMC(IS), AIS and SMC provide estimates for $\log Z$. We observe that RDSMC consistently outperforms other methods in both metrics across dimensions. Moreover, RDSMC (Proposal) exhibits consistently high weight ratio estimation bias. While RDSMC (IS) reduces this bias, it still underperforms the full RDSMC procedure, highlighting the effectiveness of intermediate resampling guided by our intermediate target distributions.
@@ 5.2 Rings and Funnel distributions
Table 1: Results on Rings and Funnel (mean $\pm$ standard error over 5 seeds). Bold indicates 95% confidence interval overlap with the best average result. RDSMC has the lowest $\log Z$ estimation bias for both targets. On Rings, RDSMC(Proposal) has the lowest radius TVD, followed by AIS, SMC, and RDSMC. On Funnel, SLIPS has the lowest sliced KSD, followed by AIS, SMC, and RDSMC.
We present results on two additional synthetic targets. Rings, introduced by grenioux2024stochastic, is a 2-dimensional distribution constructed via an inverse polar parameterization: the angular component is uniformly distributed, while the radius component follows a 4-mode Gaussian mixture. Funnel [neal2003slice], is a 10-dimensional “funnel"-shaped distribution.
For Rings, we assess sample quality using total variational distance in the radius component (Radius TVD). For Funnel, use sliced Kolmogorov-Smirnov distance (Sliced KSD) to evaluate sample quality following grenioux2024stochastic. As both targets admit tractable normalization constants, we also report the estimation bias for methods that compute them. For each method, we select the hyperparameters with lowest Raidus TVD on a heldout validation set for Rings and lowest Sliced KSD for Funnel.
As shown in Table˜1, RDSMC achieves lower or comparable estimation bias of the log-normalization constant relative to AIS and SMC on both targets. On the lower-dimensional Rings target, RDSMC (Proposal) has the lowest radius TVD, where AIS, SMC and RDSMC closely match. On the more challenging Funnel target, SLIPS achieves the lowest sliced KSD; while RDSMC performs slightly worse, it greatly outperforms its IS and Proposal variants.
@@ 5.3 Bayesian logistic regression
Finally, we evaluate inference performance on Bayesian logistic regression models using four datasets. Credit and Cancer [nishihara2014parallel] involve predicting credit risk and breast cancer recurrence, while Ionosphere [sigillito1989classification] and Sonar [gorman1988analysis] focus on classifying radar and sonar signals, respectively. The inference is performed on $60\%$ of each dataset, leaving 20% for validation and 20% for testing. Each method is tuned using the validation log-likelihood estimate. We report the test log-likelihood in Table˜2.
@@ 6 Discussion
As a nested SMC procedure, RDSMC introduces auxiliary variance. Such variance may be mitigated by enhancing autocorrelation within the MC estimation step [e.g. deligiannidis2018correlated]. In high-dimensional Gaussian mixture experiments, we observe a degree of mode collapse, an issue also seen in other samplers. Potential remedies include partial resampling [martino2016weighting], or using more informed proposals. Our method also relies on oracle metrics for hyperparameter tuning, which aligns with existing works; however, developing a more automated strategy remains an important future direction. Finally, to further improve performance, we can incorporate SNR-adapted discretization schemes and informative initializations, as explored by grenioux2024stochastic, as well as combine our approach with training-based diffusion samplers.
@@ A.1 Diffusion models
In Variance-Preserving (VP) schedule [ho2020denoising], the coefficients of the forward SDE in Eq.˜1 are given by
$f(t)=-\frac{1}{2}b(t),\quad g(t)=\sqrt{b(t)},$ (16)
where $b(t)$ is a time-dependent function
$b(t)=b_{\min}+t(b_{\max}-b_{\min}),$
@@
By Ito’s calculus (see e.g. sarkka2019applied), one can show that the forward transition density $\pi(x_{s}\mid x_{t})$ for $1\geq s>t\geq 0$ is given by
$\pi(x_{s}\mid x_{t})=\mathcal{N}\left(x_{s}\mid\frac{\alpha(s)}{\alpha(t)}x_{t},\left(1-\frac{\alpha^{2}(s)}{\alpha^{2}(t)}\right)\mathbb{I}\right),$
where
$\alpha(t)=\exp\left(-\frac{1}{2}\int_{0}^{t}b(s)ds\right),\quad\sigma(t)^{2}=1-\exp\left(-\int_{0}^{t}b(s)ds\right).$
@@
In Variance-Exploding (VE) schedule [song2020score], the forward SDE coefficients in Eq.˜1 are given by
$f(t)=0,\qquad g(t)=\sqrt{\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}},$
where $\sigma(t)$ is a non-negative and monotone increasing function with $\sigma(0)=0$. Commonly used forms of $\sigma(t)$ include polynomial functions, such as linear, quadratic or exponential schedules. When $\sigma(1)$ is large, the reference distribution can be effectively approximated as $\pi_{1}(x_{1})\approx\mathcal{N}(x_{1}\mid 0,\sigma^{2}(1))$.
Similarly to the VP case, one can derive the following parameters of the forward transition kernel with $\alpha(t)=1$ and $\sigma(t)$ is the given non-negative and monotone function.
A.2 Score identities used for diffusion MC samplers
In addition to the DSI from Eq.˜3, one may consider the following identities to form alternative MC estimates of the score, given MC samples of $\pi_{t}(x_{0}\mid x_{t})$.
@@ A.2 Score identities used for diffusion MC samplers
TSI utilizes known structural properties of the target density, such as estimated gradients, facilitating potentially more accurate and lower-variance score estimations at low-noise settings. Formally, the score function of the noisy target distribution $\pi_{t}(x)$ under the TSI framework satisfies
$\nabla_{x}\log\pi_{t}(x)$ $=\frac{1}{\alpha(t)}\int\nabla_{x}\log\pi_{t}(x_{0})\pi_{t}(x_{0}\mid x)dx_{0}.$ (17)
TSI and its related alternative have been used in several recent studies to improve score estimation accuracy [debortoli2024targetscorematching, phillips2024particle, akhound2024iterated, he2025featfreeenergyestimators].
@@
To balance the advantages of DSI and TSI across different noise regimes, MSI combines both identities through a convex interpolation. Specifically, MSI favors TSI in the low-noise regime and DSI in the high-noise regime, leveraging both strengths. Formally, MSI is a convex combination of TSI and DSI with coefficients $\frac{\alpha(t)^{2}}{\alpha(t)^{2}+\sigma(t)^{2}}$ and $\frac{\sigma(t)^{2}}{\alpha(t)^{2}+\sigma(t)^{2}}$:
$\nabla_{x}\log\pi_{t}(x)$ $=\frac{1}{\alpha(t)^{2}+\sigma(t)^{2}}\int\Big(\alpha(t)\big(x_{0}+\nabla_{x}\log\pi_{t}(x_{0})\big)-x(t)\Big)\pi_{t}(x_{0}\mid x)dx_{0}.$ (18)
he2024training leverage MSI to train an implicit generative model that produces samples in a single step, demonstrating improved sample quality and efficiency.
@@ Appendix B Method details
In this section, we discuss several details of RDSMC, including score and marginal estimation strategies in §˜B.1, the computational complexity in §˜B.2, and implementation guidelines in §˜B.3.
B.1 Score and marginal estimation
In this subsection, we present several score and marginal estimators that can be used within RDSMC, which are based on IS, AIS, SMC and rejection sampling. In our main experiments we use AIS-based estimators.
B.1.1 IS-based estimator
@@
Output: Score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ and marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
1 for $m\leftarrow 1$ to $M$ do
2   Sample ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\sim q(\cdot\mid x_{t})$
3   Compute importance weights $w^{(m)}\leftarrow\frac{\tilde{\pi}({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})\mathcal{N}(x_{t}\mid\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}},\sigma_{t}^{2}I)}{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}})}$
4
5Denote auxiliary variable ${\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\leftarrow\{{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}\}_{m=1}^{M}$
6 Compute score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\sum_{m=1}^{M}\frac{w^{(m)}}{\sum_{m^{\prime}=1}^{M}w^{(m^{\prime})}}\cdot\frac{\alpha_{t}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}^{(m)}}-x_{t}}{\sigma_{t}^{2}}$
Compute marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\leftarrow\frac{1}{M}\sum_{m=1}^{M}w^{(m)}$.
Algorithm 2 Importance Sampling (IS)-based score and marginal estimator
The IS-based estimator is summarized in Algorithm˜2, which complements the illustration in §˜3.1.
huang2023reverse consider a similar importance weighted score estimator (see their Eq 5 in Section 3.3) where the proposal $q(\cdot\mid x_{t})=\mathcal{N}(\cdot\mid x_{t}/\alpha_{t},\sigma_{t}^{2}/\alpha_{t}^{2}\mathbb{I})$ is obtained by reversing the forward noising likelihood. While the IS-based score estimator can be inefficient when $t$ is large, huang2023reverse leverages Unadjusted Langevin Algorithm (ULA) with IS estimator for initialization; And when $t$ is close to 0, they are able to quickly obtain rough score estimates via the IS approach.
B.1.2 AIS and SMC-based estimator